Hi the simultaneous equation is
2v=31-3w
3v+w=64

Found 2 solutions by ankor@dixie-net.com, Alan3354:
Answer by ankor@dixie-net.com(22740)

the simultaneous equation is
2v = 31 - 3w rewrite to 2v + 3w = 31
3v + w = 64 multiply by 3
then we have
9v + 3w = 192
2v + 3w = 31
---------------subtraction eliminates w, find v
7v + 0 = 161
v = 161/7
v = 23
:
Find w using the equation 3v + w = 64, replace v with 23
3(23) + w = 64
69 + w = 64
w = 64 - 69
w = -5
:
:
Check these solutions in the equation 2v = 31 - 3w
2(23) = 31 - 3(-5)
46 = 31 + 15

Answer by Alan3354(69443)

2v=31-3w
3v+w=64 --> w = -3v+64
-----------
Sub for w in the 1st equation.